Allan bought some 5 cent stamps and some 20 cent stamps. The entire purchase was $7 he bought three times as many f cent stamps as 20 cent stamps how many of each stamps did he buy?
Answer by oberobic(2304)

x = number of five-cent stamps
5x = value of the five-cent stamps in cents
y = number of twenty-cent stamps
20y = value of the twenty-cent stamps in cents
.
5x +20y = 700 cents
.
three times as many 5-cent stamps as 20-cent stamps
x = 3*y
.
substitute
.
5(3y) +20y = 700
15y + 20y = 700
35y = 700
y = 20
.
x =3y = 3(20) =60
.
check the value
.
5(60) = 300 cents
20(20) = 400 cents
300+400 = 700 cents
correct
.
Answer: He bought 20 twenty-cent stamps and 60 five-cent stamps.